Hello,

I need help solving this word problem using a system of equations:

A boat travels 60 miles downstream in the same time it takes to go 36 miles upstream. The speed of the boat in still water is 15 mph greater than the speed of the current. Find the speed of the current.

Thank you

Answer by scott8148(5891)

let x=(boat in still water); let c=current ... so x=c+15 and c=x-15

t=d/r ... the times are the same so 60/(x+c)=36/(x-c) ... cross multiplying gives 60(x-c)=36(x+c) ... 60x-60c=36x+36c

substitute from the first line to find x or c; then use an equation from the first line to find the other one